# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ Appendix A Theoretical Analysis
A.1 Equivariance of NLSFs
We start with two lemmas that characterize the graph functional shifts. The lemmas directly follow the fact that two normal operators commute iff the projections upon their eigenspaces commute.
@@ A.1 Equivariance of NLSFs
Denote by $\mathcal{R}$ the span of $\bigcup_{j=1}^{J}\mathbf{P}_{j}{\mathbb{R}}^{N}$. Denote by $\mathcal{R}^{\perp}$ the orthonormal complement of $\mathcal{R}$. Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{J}\mathbf{P}_{j}{\mathbb{R}}^{N}\right)\times\mathcal{R}^{\perp},$
where $\prod$ and $\times$ denote direct products of linear spaces. Denote by $\oplus$ the direct sum of operators.
Lemma A.1.
$\mathbf{U}$ is in $\mathcal{U}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=(\bigoplus_{j=1}^{J}\mathbf{U}_{j})\oplus\mathbf{U}^{\perp}$, where $\mathbf{U}_{j}$ is a unitary operator in $\mathbf{P}_{j}{\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}$.
@@
Denote by $\mathcal{R}_{g}$ the span of $\bigcup_{j=1}^{K}g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$. Denote by $\mathcal{R}^{\perp}_{g}$ the orthonormal complement of $\mathcal{R}_{g}$ Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{K}g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}\right)\times\mathcal{R}^{\perp}_{g}.$
Lemma A.2.
$\mathbf{U}$ is in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=(\bigoplus_{j=1}^{K}\mathbf{U}_{j})\oplus\mathbf{U}^{\perp}_{g}$, where $\mathbf{U}_{j}$ is a unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}_{g}$.
A.1.1 Proof of Propositions 4.1
@@
Index NLSFs in Eq. (4) are equivariant to the graph functional shift $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@
We start with Index-NLSFs. Consider the Index NLSFs defined as in Eq. (4)
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
@@
Consider $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$ and apply it to the graph signal $\mathbf{X}$, the Index NLSFs with the transformed input is given by
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
Since $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$, it commutes with $\mathbf{P}_{j}$ for all $j$, i.e., $\mathbf{U}\mathbf{P}_{j}=\mathbf{P}_{j}\mathbf{U}$. Using this commutation property, we can rewrite the norm and the projections:
@@
Substitute these expressions back into the definition of the Index NLSFs:
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
Notice that $\mathbf{U}$ appears linearly in the numerator of the fraction. We can factor it out:
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\mathbf{U}\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
The expression inside the summation is exactly the original Index NLSFs applied to
@@
A.2 Expressivity and Universal Approximation
@@
A.2.1 Proof of Lemma 4.1 - Linear Node-level NLSF Exhaust the Linear Operators that Commute with Functional Shifts
Lemma 4.1
Suppose that $\sum_{j}\mathbf{P}_{j}=\sum_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$. A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commute with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
@@
Consider a setting where a single graph is given as a fixed domain, and the data distribution consists of many signals defined on this graph. For simplicity, we restrict the analysis to the case of 1D signals ($d=1$). The extension to a general dimension $d$ is natural.
By Lemma A.2, any unitary operator $\mathbf{U}$ in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ has the form
$\mathbf{U}=\left(\bigoplus_{j=1}^{K}\mathbf{U}_{j}\right)\oplus\mathbf{U}^{\perp},$
where $\mathbf{U}_{j}$ is a unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}_{g}$.
Hence, since the unitary representation $\mathbf{U}\mapsto\mathbf{U}$ of the group of unitary operators in ${\mathbb{R}}^{N}$ is irreducible, by Schur’s lemma [Serre et al., 1977] any linear operator $\mathbf{B}$ that commutes with all operators of $\mathcal{U}^{g}_{{\bm{\Delta}}}$ must be a scalar times the identity when projected to $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$ and to $\mathcal{R}_{g}^{\perp}$. Namely, $\mathbf{B}$ has the form
$\mathbf{B}=\left(\bigoplus_{j=1}^{K}(b_{j}\mathbf{I}_{j})\right)\oplus(b^{\perp})\mathbf{I}^{\perp},$
where $\mathbf{I}_{j}$ is the identity operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, $\mathbf{I}^{\perp}$ is the identity operator in $\mathcal{R}^{\perp}_{g}$, and $b_{j}$ and $b^{\perp}$ are scalars. This means that linear node-level NLSFs exhaust the space of linear operators that commute with $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@
NLSF with High Frequency Content.
We note that the condition $\sum_{j}\mathbf{P}_{j}=\sum_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$ can be omitted from Lemma 4.1 if we extend the definition of NLSF by adding one more term in the spectrum as follows. In the spectral index case, concatenate to the analysis operator the coefficient $\left\lVert\mathbf{X}-\sum_{j}\mathbf{P}_{j}\mathbf{X}\right\rVert$, and in the value case, concatenate $\left\lVert\mathbf{X}-\sum_{j}g_{j}({\bm{\Delta}})\mathbf{X}\right\rVert$. Extend accordingly the definition of synthesis and of NLSFs to also depend on this term. We do not include this spectral term in practice as it does not lead to improved performance.
A.2.2 Node-Level Universal Approximation
The above analysis motivates the following construction of a metric on ${\mathbb{R}}^{N}$. Given the filter bank $\{g_{j}\}_{j=1}^{K}$, on graph with $d$-dimensional signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined to be
@@ A.2.2 Node-Level Universal Approximation
Now, since MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{K\times d^{\prime}}$ (by the universal approximation theorem), and by the fact that $\mathcal{A}_{{\bm{\Delta}}}/{\rm dist}_{{\bm{\Delta}}}$ is an isometry, node-level NLSFs based on MLPs in the spectral domain can approximate any continuous function from signals with $d$ features to signal with $d^{\prime}$ features $(\mathcal{N}_{{\bm{\Delta}}})^{d}\rightarrow(\mathcal{N}_{{\bm{\Delta}}})^{d^{\prime}}$.
A.2.3 Graph-Level Universal approximation
The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two graphs with Laplacians and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$, their distance is defined to be
@@ A.2.3 Graph-Level Universal approximation
By the isometry property, $\mathcal{G}$ inherits any approximation property from ${\mathbb{R}}^{K\times d}$. For example, since MLPs can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{d^{\prime}}$, the space of NLSFs based on MLPs $\Psi$ has a universality property: any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by a NLSF based on an MLP.
A.2.4 Graph-Level Expressivity of Pooling-NLSFs
@@
A.3 Stable Invertibility of Synthesis

### v2

@@ Appendix D Theoretical Analysis
D.1 Equivariance of NLSFs
We start with two simple lemmas that characterize the graph functional shifts. The lemmas directly follow the fact that two normal operators commute iff the projections upon their eigenspaces commute.
@@ D.1 Equivariance of NLSFs
Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{J+1}\mathbf{P}_{j}{\mathbb{R}}^{N}\right),$
where $\prod$ denotes direct products of linear spaces and the $(J+1)$-th filter is the orthogonal complement, given by $\mathbf{P}_{J+1}=\mathbf{I}-\sum_{j=1}^{J}\mathbf{P}_{j}$. Denote by $\oplus$ the direct sum of operators.
Lemma D.1.
$\mathbf{U}$ is in $\mathcal{U}^{J}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=\left(\bigoplus_{j=1}^{J+1}\mathbf{U}_{j}\right)$, where $\mathbf{U}_{j}$ is a unitary operator in $\mathbf{P}_{j}{\mathbb{R}}^{N}$.
Proof.
Let $\mathbf{U}\in\mathcal{U}^{J}_{{\bm{\Delta}}}$. Since $\mathbf{U}$ commute with $\mathbf{P}_{j}$ for $j=1,\ldots,J$, and with $\mathbf{I}$, it also commutes with $\mathbf{P}_{J+1}=\mathbf{I}-\sum_{j=1}^{J}\mathbf{P}_{j}$. Therefore, we can write
$\mathbf{U}=\sum_{j=1}^{J+1}\mathbf{P}_{j}\mathbf{U}=\sum_{j=1}^{J+1}\mathbf{P}_{j}^{2}\mathbf{U}=\sum_{j=1}^{J+1}\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}.$
Now, when restricting $\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}$ to an operator in $\mathbf{P}_{j}{\mathbb{R}}^{N}$, it is unitary. Indeed, for every $\mathbf{v}=\mathbf{P}_{j}\mathbf{v}\in\mathbf{P}_{j}{\mathbb{R}}^{N}$ and $\mathbf{u}=\mathbf{P}_{j}\mathbf{u}\in\mathbf{P}_{j}{\mathbb{R}}^{N}$, since $\mathbf{U}$ is unitary and $\mathbf{P}_{j}$ is self-adjoint and satisfies $\mathbf{P}_{j}^{2}=\mathbf{P}_{j}$, we have
$\left\langle\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}\mathbf{v},\mathbf{u}\right\rangle=\left\langle\mathbf{v},\mathbf{P}_{j}\mathbf{U}^{-1}\mathbf{P}_{j}\mathbf{u}\right\rangle,$
and $\mathbf{P}_{j}\mathbf{U}^{-1}\mathbf{P}_{j}$ is the inverse of $\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}$ in $\mathbf{P}_{j}{\mathbb{R}}^{N}$, since for every $\mathbf{v}\in\mathbf{P}_{j}{\mathbb{R}}^{N}$
$\mathbf{P}_{j}\mathbf{U}^{-1}\mathbf{P}_{j}\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}\mathbf{v}=\mathbf{P}_{j}\mathbf{U}^{-1}\mathbf{U}\mathbf{P}_{j}\mathbf{v}=\mathbf{P}_{j}\mathbf{v}=\mathbf{v},$
and similarly $\mathbf{P}_{j}\mathbf{U}\mathbf{P}_{j}\mathbf{P}_{j}\mathbf{U}^{-1}\mathbf{P}_{j}\mathbf{v}=\mathbf{v}$. Here, an invertible normal operator commutes with an orthogonal projection if and only if its inverse does.
The other direction is trivial. ∎
@@
Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{K+1}g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}\right),$
where the $(K+1)$-th filter is the orthogonal complement, given by $g_{K+1}({\bm{\Delta}})=\mathbf{I}-\sum_{j=1}^{K}g_{j}(\mathbf{\Delta})$.
Lemma D.2.
$\mathbf{U}$ is in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=\left(\bigoplus_{j=1}^{K+1}\mathbf{U}_{j}\right)$, where $\mathbf{U}_{j}$ is a unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$.
The proof is analogous to the proof of Lemma D.1.
D.1.1 Proof of Propositions 4.1
@@
Index NLSFs in Eq. (4) are equivariant to the graph functional shifts $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@
We start with Index-NLSFs. Consider the Index NLSF defined as in Eq. (4)
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J+1}.$
@@
Consider $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$ and apply it to the graph signal $\mathbf{X}$. The Index NLSF with the transformed input is given by
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\left[\frac{\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J+1}.$
Since $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$, it commutes with $\mathbf{P}_{j}$ for all $j$, i.e., $\mathbf{U}\mathbf{P}_{j}=\mathbf{P}_{j}\mathbf{U}$. Using this commutation property, we can rewrite the norm and the projections as
@@
Substituting these expressions back into the definition of the Index NLSFs gives
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\left[\frac{\mathbf{U}\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J+1}.$
Notice that $\mathbf{U}$ appears linearly in the numerator of the fraction. Hence, we can factor it out by
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\mathbf{U}\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\left[\Psi_{\text{ind}}\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\right]_{i=1}^{J+1}.$
The expression inside the summation is exactly the original Index NLSFs applied to ${\mathbf{X}}$, so
@@
D.2 Expressivity and Universal Approximation
@@
D.2.1 Proof of Lemma 4.1 - Linear Node-level NLSF Exhaust the Linear Operators that Commute with Functional Shifts
Lemma 4.1.
A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commute with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
@@
For simplicity, we restrict the analysis to the case of 1D signals ($d=1$). The extension to a general dimension $d$ is natural.
By Lemma D.2, the unitary operators $\mathbf{U}$ in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ are exhausted by the operators of the form
$\mathbf{U}=\left(\bigoplus_{j=1}^{K+1}\mathbf{U}_{j}\right),$
where $\mathbf{U}_{j}$ is any unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$. Hence, since the unitary representation $\mathbf{T}\mapsto\mathbf{T}$ of the group of unitary operators in ${\mathbb{R}}^{m}$ (for any $m\in\mathbb{N}$) is irreducible, by Schur’s lemma [85] any linear operator $\mathbf{B}$ that commutes with all operators of $\mathcal{U}^{g}_{{\bm{\Delta}}}$ must be a scalar times the identity when projected to $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$. Namely, $\mathbf{B}$ has the form
$\mathbf{B}=\left(\bigoplus_{j=1}^{K+1}(b_{j}\mathbf{I}_{j})\right),$
where $\mathbf{I}_{j}$ is the identity operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$. This means that linear node-level NLSFs exhaust the space of linear operators that commute with $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
@@
D.2.2 Node-Level Universal Approximation
The above analysis motivates the following construction of a metric on ${\mathbb{R}}^{N}$. Given the filter bank $\{g_{j}\}_{j=1}^{K+1}$, on graph with $d$-dimensional signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{(K+1)\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined to be
@@ D.2.2 Node-Level Universal Approximation
Now, since MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{(K+1)\times d}\rightarrow{\mathbb{R}}^{(K+1)\times d^{\prime}}$ (by the universal approximation theorem), and by the fact that $\mathcal{A}_{{\bm{\Delta}}}/{\rm dist}_{{\bm{\Delta}}}$ is an isometry, node-level NLSFs based on MLPs in the spectral domain can approximate any continuous function from signals with $d$ features to signal with $d^{\prime}$ features $(\mathcal{N}_{{\bm{\Delta}}})^{d}\rightarrow(\mathcal{N}_{{\bm{\Delta}}})^{d^{\prime}}$.
D.2.3 Graph-Level Universal approximation
The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{(K+1)\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two graphs with Laplacians and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$, their distance is defined to be
@@ D.2.3 Graph-Level Universal approximation
By the isometry property, $\mathcal{G}$ inherits any approximation property from ${\mathbb{R}}^{(K+1)\times d}$. For example, since MLPs can approximate any continuous function ${\mathbb{R}}^{(K+1)\times d}\rightarrow{\mathbb{R}}^{d^{\prime}}$, the space of NLSFs based on MLPs $\Psi$ has a universality property: any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by a NLSF based on an MLP.
D.2.4 Graph-Level Expressivity of Pooling-NLSFs
@@
D.2.5 Discussion on Graph-Level Expressivity of NLSFs
We offer additional discussion on the expressivity of graph-level NLSFs. Note that graph-level NLSFs are bounded by the expressive power of MPNNs with random positional encodings. For example, in [81], the authors showed that random features improve GNN expressivity, distinguishing certain structures that deterministic GNNs cannot. The Lanczos algorithm for computing the leading $J$ eigenvectors can be viewed as a message-passing algorithm with a randomly initialized $J$-channel signal.
The example in App. D.2.4 can be written as a standard linear graph spectral filter, followed by the pooling (readout) function. This shows that graph-level NLSFs (without synthesis and pooling) are not more expressive than standard spectral GNNs with pooling. In the other direction, there are no graph-signals that can be separated by a graph-level NLSF but not by a spectral GNN. Given two graph-signals that an NLSF can separate, we can build a specific standard spectral GNN that gives the same output as the NLSF specifically for the two given graph-signals. However, several considerations should be noted:
  1. 1.
    the feature dimension of this GNN would is as large as the combined number of eigenvalues of the two graphs times the number of features,
  2. 2.
    this GNN is designed to fit these two specific graphs, and will not work for other graphs, and,
  3. 3.
    if implemented via polynomial filters, the polynomial order would have to be the combined number of eigenvalues of the two graphs, which makes it very unstable.
Let us explain the architecture next. Given the two graphs $(G_{1},f_{1})$ with $N_{1}$ vertices and $(G_{2},f_{2})$ with $N_{2}$ vertices, with node features of dimension $D$, define band-pass filters that separate the combined set of eigenvalues of the two given graphs. Concatenate these filters to build a single multi-channel filter. Namely, this filter maps the signal to the sequence of band-pass filtered signal – each band at a different channel, for a total of $D(N_{1}+N_{2})$ channels. If the band-pass filters are based on polynomials, each polynomial would be chosen to be zero on all eigenvalues except for one. Apply $L_{2}$-norm pooling on each channel to obtain a single feature of dimension $D(N_{1}+N_{2})$. This gives the sequence of norms of the signal at the bands, where, for the two given graphs, the two pooled signals of the two graphs are supported on disjoint sets of indices. Hence, the linear spectral filter contains the frequency information of the NLSF. Then, one can apply the MLP of the NLSF to the channels corresponding to the first graph, and to the channels corresponding to the second graph. This means that the linear spectral GNN gives the exact same outputs on $(G_{1},f_{1})$ and $(G_{2},f_{2})$ as the NLSF.
Regarding pooling-NLSF, we do not offer an answer to the question whether standard spectral GNNs are more powerful (assuming an unlimited budget) than pooling-NLSFs, or vice-versa. More practically meaningful questions would be:
  1. 1.
    Compare the expressivity of standard spectral GNNs to NLSFs for a given budget of parameters.
  2. 2.
    How many graph-signal can a single spectral GNN separate, vs a single NLSF, of the same budget.
We leave these questions as open problems for future research. We note that, in practice, NLSFs with the same budget as standard spectral GNNs perform better.
D.3 Stable Invertibility of Synthesis